Project

General

Profile

NDOS Commissioning Meeting 2011/05/05

Attendance: Jon, Susan, Jan, Athans, Carl, Mark Messier, Andrew, Denis, Peter, Jarek, Chad, Mat, Kurt

Agenda:

  1. NDOS Status and Channel Recovery Plans (Jon, Mat, Denis)
  2. DAQ/Electronics Update & Plans (Peter, Andrew, Kurt)
  3. DCS Update & Plans (Athans, Chad)
  4. DataCheck update (UMN)
  5. AOB (All)

NDOS Status and Channel Recovery Plans (Jon, Mat, Denis)

  • dcm 4-2? Pedestal scans looked reasonable, but hasn't been in the readout...
  • cap-board installation is complete on entire detector. Will need to determine success rate. Denis believes the success rate is better than previously reported. Cooling does seem to make the noise worse on some (~15-20%) APDs. Denis will continue tests on Monday, and should have a summary report next Thursday.
  • Mat just finished diblock 2, so IPND is done. Top-off will continue tomorrow. Will probably need Monday to finish it off.
  • Mat has summarized the current status of NDOS installation/channel counts in doc-dB
  • Leon reported that 20 APDs will be sent to FNAL as soon as he's done commissioning the APD test stand at CalTech. This should be in ~2 weeks.

DAQ/Electronics Update & Plans (Peter, Andrew, Kurt)

  • programs were spontaneously dying on daq-master today. Andrew: possibly a networking problem.
  • new version of the message facility was tested successfully yesterday. We're back to using the older version today, plan is to move permanently to the new version early next week.
  • the DCMs will be renamed to their logical names on the detector. This involves more book-keeping behind the scenes, but will simplify life for many others. This will happen next week.
  • DAQ stress-tests are planned for tomorrow. This will involve scaling trigger/data rates up. Will start with pulsers first.

DCS Update & Plans (Athans, Chad)

  • Chad: reorganized wiring at NDOS; patch planes will simplify things. We do not have a transmitting flowmeter, so this will have to be added to the shifter walk-through checklist for the short term. Chad will work on ordering flowmeters, but getting someone (or two) to do the installation is a bigger challenge.
  • Athans:
    • has asked the DAQ group for help to set up the test-stand for long tests of the DCS+DAQ configuration. Will start tests tomorrow and will run over the weekend.
    • has started putting together a Requirements document for DCS. See doc-dB 5946. Help is requested to add content to the document.

DataCheck update

AOB (All)